d^2y/dx^2–cotxdy/dx–sinx^2 y=cosx–cos^3x
1
Expert's answer
2020-02-14T10:13:05-0500

SOLVE d 2 y/d x 2-cot x d y/d x-sin 2 x y=cos x-cos 3 x

Solution

P=- cot x Q=- sin 2 x R=cos x-cos 3 x

choose , z such that

(dz/d x) 2 = sin 2 x

taking square root on both sides we get

dz/d x=sin x

∫ dz/d x =∫ sin x d x

z= - cos x

z =- cos x

derivative with respect x

dz/d x= sin x

2 z/d x 2= cos x

1= (d 2 z/d x 2+ P dz/d x )/(dz/d x) 2

1=cos x-cot x sin x/sin 2 x

1= {cos x-(cos x/sin x)sin x}/sin 2 x

1=cos x-cos x/sin 2 x

1=0

1=Q/(dz/d x)2

1= - sin 2 x/sin 2 x

1 = Q/(dz/d x) 2

1 = (cos x-cos 3 x)/sin 2 x

1= -z-(-z 3)/{1-(-z)2}

1= -z+z 3/1-z 2

1= - z(-z 2+1)/(1-z 2)

1 = -z

Reduced Equation is

2 y/d z 2-y= -z .......>(1)

Complementary solution:-

Auxiliary Equation is

2 -1=0

m=±1

y c(x)=C 1 e z+C 2 e -z

Particular Solution:-

y p(x)= -A z (2)

derivative w.r.t z

d y/dz= -A

again derivative

2 y/d z 2=0

by putting the value in equation no 1

0+A z=-z

A=-1

putting in equation no 2

y p(x)= z

y(x)=y c(x)+y p(x)

y(x)=C 1 e z+C 2 e -z +z

Hence The complete solution is

y(x)=C 1 e z+ C 2 e -z+z

y(x)=C 1 e -cos x+ C 2 e cos x -cos x Answer